# Defactify 4.0: Multimodal Fact-Checking and AI-Generated Image Detection, March 2025, Philadelphia, Pennsylvania, USA

### v1

@@ 1 Introduction
By providing a large-scale, representative benchmark, our dataset aims to advance research in synthetic media detection and foster the development of scalable countermeasures against AI-enabled disinformation. Building upon the foundations laid by initiatives such as the Defactify workshop series [defactify], this work bridges the gap between academic inquiry and practical implementation, offering a valuable resource for researchers, policymakers, and industry stakeholders committed to safeguarding the integrity of digital information ecosystems. This work complements parallel efforts addressing AI-generated text detection [roy2025comprehensive].
@@ 2.1 AI-Generated Image Datasets
  • GenImage: Zhu2023GenImage built a million-scale benchmark with AI-generated and real image pairs. While large in scale, the dataset mainly features older generators and lacks fine-grained model labels, limiting its use for studying modern diffusion models.
@@ 2.2 Detection Methods
Our dataset provides caption-aligned real and synthetic images from five modern generators (Stable Diffusion 3 [esser2024scalingrectifiedflowtransformers], Stable Diffusion 2.1 [rombach2022high], SDXL [podell2023sdxlimprovinglatentdiffusion], DALL-E 3 [ramesh2021zero], MidJourney v6 [holzmidjourney]), with model attribution labels and systematic perturbations for robustness evaluation.
@@ 3.1 Image generation and Annotation
Our dataset is built upon the MS COCO dataset [lin2014microsoft], which provides high–quality real images paired with human-written captions. We randomly sample 16k image-caption pairs from the MS COCO dataset. Each caption is used as a textual seed for generating synthetic images using five image-generation models - Stable Diffusion 3 (SD3 ) [esser2024scalingrectifiedflowtransformers], Stable Diffusion 2.1 (SD 2.1) [rombach2022high], SDXL [podell2023sdxlimprovinglatentdiffusion], DALL-E 3 [ramesh2021zero], MidJourney v6 [holzmidjourney]. For every caption, each model produced one synthetic image, resulting in a multi-source collection of AI-generated visual samples.
@@ 3.3 Data Structure
  • label_1 — Binary label indicating whether the image is real (0) or AI-generated (1).
  • label_2 — Categorical label indicating the specific generative model (SD 3, SDXL, SD 2.1, DALL-E 3, or MidJourney 6).
@@ 3.4 Data Analysis
4 CT2 - AI Generated Image Detection Tasks
@@ 5 Baseline
To establish a baseline, we train a ResNet-50 classifier using image representations in the frequency domain. These frequency domain representations are generated by applying a pre-processing strategy inspired by the methodology presented in [Corvi_2023_CVPR]. This transformation captures global frequency characteristics that help reveal subtle artifacts often present in synthetic images.
@@ 6 Results
Table 3: Baseline results for both tasks in the shared task.
@@ 7 Conclusion
In this paper, we release a large-scale dataset for AI-generated image detection comprising 96,000 real and synthetic image-caption pairs. A key feature of our dataset is semantic alignment- all synthetic images are generated from the same captions as their real counterparts, enabling controlled studies that separate content bias from generation artifacts.

### v2

# DeFactify 4.0: Fourth workshop on Multimodal Fact-Checking and Hate Speech Detection, March 2025, Philadelphia, Pennsylvania, USA

@@ 1 Introduction
By providing a large-scale, representative benchmark, our dataset aims to advance research in synthetic media detection and foster the development of scalable countermeasures against AI-enabled disinformation. Building upon the foundations laid by initiatives such as the Defactify workshop series [defactify], this work bridges the gap between academic inquiry and practical implementation, offering a valuable resource for researchers, policymakers, and industry stakeholders committed to safeguarding the integrity of digital information ecosystems. This work complements parallel efforts addressing AI-generated text detection [roy-2025-defactify-dataset-text].
@@ 2.1 AI-Generated Image Datasets
  • GenImage: Zhu2023GenImage built a million-scale benchmark with AI-generated and real image pairs. While large in scale, the mainly features older generators and lacks fine-grained model labels, limiting its use for studying modern diffusion models.
@@ 2.2 Detection Methods
Beyond model-specific artifact detection, benchmark-oriented evaluations such as the Visual Counter Turing Test (VCT2) [imanpour2025visual] have been proposed to systematically assess detection robustness across generative models, introducing difficulty-aware metrics like the Visual AI Index (VAI).
Our dataset provides caption-aligned real and synthetic images from five modern generators (Stable Diffusion 3 [esser2024scalingrectifiedflowtransformers], Stable Diffusion 2.1 [rombach2022high], SDXL [podell2023sdxlimprovinglatentdiffusion], DALL-E 3 [BetkerImprovingIG], MidJourney v6 [holzmidjourney]), with model attribution labels and systematic perturbations for robustness evaluation.
@@ 3.1 Image generation and Annotation
Our dataset is built upon the MS COCO dataset [lin2014microsoft], which provides high–quality real images paired with human-written captions. We randomly sample 16k image-caption pairs from the MS COCO dataset. Each caption is used as a textual seed for generating synthetic images using five image-generation models - Stable Diffusion 3 (SD3 ) [esser2024scalingrectifiedflowtransformers], Stable Diffusion 2.1 (SD 2.1) [rombach2022high], SDXL [podell2023sdxlimprovinglatentdiffusion], DALL-E 3 [BetkerImprovingIG], MidJourney v6 [holzmidjourney]. For every caption, each model produced one synthetic image, resulting in a multi-source collection of AI-generated visual samples.
@@ 3.3 Data Structure
  • label_A — Binary label indicating whether the image is real (0) or AI-generated (1).
  • label_B — Categorical label indicating whether an image is real (0) or generated by a specific generative model: Stable Diffusion 2.1 (1), Stable Diffusion XL (2), Stable Diffusion 3 (3), DALL-E 3 (4), or Midjourney 6 (5).
@@ 3.4 Data Analysis
4 Counter Turing Test (CT2) - AI Generated Image Detection Tasks
@@ 5 Baseline
To establish a baseline, we train a ResNet-50 classifier using image representations in the frequency domain. These frequency domain representations are generated by applying a preprocessing strategy inspired by the methodology presented in [Corvi_2023_CVPR]. This transformation captures global frequency characteristics that help reveal subtle artifacts often present in synthetic images.
@@ 6 Results
Table 3: Baseline results for both tasks.
@@ 7 Conclusion
In this paper, we release a large-scale dataset for AI-generated image detection comprising 96,000 real and synthetic image-caption pairs. A key feature of our dataset is semantic alignment, all synthetic images are generated from the same captions as their real counterparts, enabling controlled studies that separate content bias from generation artifacts.